OpenStudy (brittany):

Simplify and show work 5+√3/4+√3

OpenStudy (anonymous):

\[((5+\sqrt{3})\times (4-\sqrt{3}))/(4+\sqrt{3})\times (4-\sqrt{3})\]\[(17 -\sqrt{3})/13\]

OpenStudy (brittany):

How do you multiply them? like which parts are multiplied befor u get that ansseer